OpenStudy (mitu12):

(x^9)^0 (x^7)^2 SIMPLIFY

OpenStudy (cwrw238):

anything to the power 0 = 1

OpenStudy (mitu12):

so it would be x^9

OpenStudy (cwrw238):

no (x^a)^b = x^ab

OpenStudy (cwrw238):

(x^7)^2 = x^(2*7)

OpenStudy (mitu12):

x^14

OpenStudy (cwrw238):

yes

OpenStudy (mitu12):

could you show me the work please

OpenStudy (cwrw238):

(x^9)^0 (x^7)^2 = 1 * (x^(7*2) = x^14

OpenStudy (mitu12):

ok thank you

OpenStudy (cwrw238):

yw